### v1

# Prompt Mixing in Diffusion Models using the Black Scholes Algorithm

Abstract
We introduce a novel approach for prompt mixing, aiming to generate images at the intersection of multiple text prompts using pre-trained text-to-image diffusion models. At each time step during diffusion denoising, our algorithm forecasts predictions w.r.t. the generated image and makes informed text conditioning decisions. To do so, we leverage the connection between diffusion models (rooted in non-equilibrium thermodynamics) and the Black-Scholes model for pricing options in Finance, and draw analogies between the variables in both contexts to derive an appropriate algorithm for prompt mixing using the Black Scholes model. Specifically, the parallels between diffusion models and the Black-Scholes model enable us to leverage properties related to the dynamics of the Markovian model derived in the Black-Scholes algorithm. Our prompt-mixing algorithm is data-efficient, meaning it does not need additional training. Furthermore, it operates without human intervention or hyperparameter tuning. We highlight the benefits of our approach by comparing it qualitatively and quantitatively to other prompt mixing techniques, including linear interpolation, alternating prompts, step-wise prompt switching, and CLIP-guided prompt selection across various scenarios such as single object per text prompt, multiple objects per text prompt and objects against backgrounds. Code is available at https://github.com/divyakraman/BlackScholesDiffusion2024.
Figure 1: Our method’s results (Black Scholes, last column) are presented alongside comparisons to prior work. Vanilla stable diffusion (SD) struggles to capture clear characteristics of individual text prompts (notably missing distinct features such as those of the parrot, cat, dog/penguin, and sunset/penguin). Linear interpolation performs poorly due to non-linear manifolds. Alternating sampling and step-wise switching yield low-quality results with artifacts, primarily because they lack intelligent prompt selection during denoising steps (missing characteristics of pizza, artifacts in cat/muffin and dog/penguin mixing, sunset/northern lights not well captured). CLIP-min exhibits bias issues by not modeling diffusion denoising dynamics and prompt selection effectively, which hinders fore-sighted decision making, the generated images are biased towards one of the text prompts. In contrast, our Black-Scholes model generates realistic images that meticulously balance and preserve the characteristics of each individual text prompt. The images are from set 1, set 2, set 3 and set 4 (Refer Section 5.1) respectively.
1 Introduction
Text-to-image diffusion models [13, 14] such as Imagen [33] by Google, Firefly [2] by Adobe, DALLE [7] by Microsoft, and Stable Diffusion [31, 11] by Stability AI can produce photorealistic images from textual descriptions and are typically trained on vast amounts of data using substantial computational resources. The latest versions of these models (as of May 2024) have seen significant improvements in comparison to their initial versions, attributed to various factors, such as increased training data, enhanced text understanding networks, improved image generation networks, and optimized strategies. Data efficient solutions [36], leveraging pre-trained diffusion models, can enhance results without requiring additional data collection. Prompt engineering [37] is one effective approach, involving techniques such as rephrasing prompts to improve model generalization. Another method is using large language models (LLMs) [23, 38] to parse complex prompts and determine useful priors for image generation. Attention maps [9] and prompt mixing techniques [27, 5] also contribute to better results.
Prompt mixing becomes particularly valuable when aiming to generate images that blend different text-image concepts i.e. when aiming to generate images at the intersection [27, 30, 28] of different text-image concepts. For instance, consider generating an image resembling a hybrid of a pink cat and a dog. Current diffusion models [2, 33, 7, 11] excel at straightforwardly combining individual text prompts to create visually appealing combinations. However, these capabilities are often restricted to advanced model versions (version 3) and rely on large-data techniques.
Data efficient prompt mixing techniques can enhance image generation without additional data. While linear interpolation [17] of text embeddings from individual prompts is a simple approach, it may not be optimal due to the highly non-linear nature of the text-image manifold and potential bias issues. Alternatives include switching between prompts during diffusion denoising, either by alternating [18] or using a step-wise technique [27]. However, these methods often require human involvement and careful prompt engineering for optimal results.
The question arises: what’s the best way to switch between text prompts for prompt-mixing? An automated approach considers the model’s varying capabilities with different prompts. While attention maps [15, 9] and layout guidance![42, 10] methods excel at guiding the model toward distinct scene entities, they may not be as effective for blending concepts within the same entity. To address this, during each step of diffusion denoising, the network should automatically focus on the text prompt that corresponds to the image’s deficiencies.
Main contributions.
In this paper, we introduce a novel approach to prompt mixing by integrating concepts from economics and finance. Our objective is to leverage pretrained diffusion models for prompt-mixing to generate images at the intersection of various text-image manifolds. During each step of the diffusion denoising process, our method dynamically conditions on the text prompt that requires the highest level of attention. To achieve this, our algorithm assesses the ‘cost’ associated with each relevant text prompt and selects the optimal conditions for the subsequent step of image generation based on the ‘cost’ that requires maximum optimization.
Key Insight: The inspiration for our approach lies in non-equilibrium thermodynamics [35], which is at the core of the development of diffusion models. These models share a conceptual foundation with the Black Scholes model[26], a Nobel Prize-winning mathematical framework extensively employed in financial markets for pricing European call options[8]. The denoising steps within a diffusion model, aimed at image generation, constitute a Markovian time-series. In our analogy, the image itself represents a valuable ‘stock’ or ‘asset’ that we aim to ‘purchase’ (or generate) at the most favorable ‘cost’ (or alignment with the text prompts). During each step of the diffusion denoising process, we extrapolate the generate latents to compute the ‘stock’ prices at the corresponding time-step, and use the properties of diffusion models to compute the various variables involved in the Black Scholes algorithm. Consequently, we leverage the Black Scholes model to predict a score for each text prompt. This score serves as an indicator of how the image should be conditioned in the subsequent timestep. By adopting this approach, our model dynamically focuses on aspects that require attention, ultimately generating an image that optimally satisfies all relevant text prompts.
We perform experiments on prompts with varying complexities to assess our method’s ability to seamlessly blend different objects and backgrounds across scenarios. Through qualitative and quantitative comparisons against several baselines-such as linear interpolation, prompt switching, and CLIP score-we highlight the superiority of our approach.
2 Related work
Prompt mixing.
Prompt mixing is a technique where different text prompts are used at various steps during the diffusion denoising process. Patashnik et al. [27] follow a step-wise approach, while Aerial Diffusion [18] alternates between prompts to create semantically consistent aerial-view images. Tools such as Image Mixer Diffusion [28] and CLIP Guided Image Mixing [30] follow similar techniques of prompt switching to blend multiple text prompts. However, these methods often require complex hyperparameter tuning.
Mixing step.
The mixing time or mixing step [22] of a Markov chain refers to the duration it takes for the chain to reach its steady-state distribution. Mixing time has found applications in image editing [46] and synthesizing out-of-distribution (OOD) images [45]. When switching between prompts using mixing time, a step-wise approach is followed, akin to the work by Patashnik et al. [27]. The advantage of mixing-time approaches over the method proposed by Patashnik et al. [27] is that it mathematically determines the optimal switching time without requiring complex hyperparameter tuning. To approximate the mixing time, it estimates the radius of the latent space. However, there are a few limitations to these methods. Firstly, most of these approaches are applied to diffusion models trained on relatively small, problem-specific datasets. Secondly, identifying clear boundaries on large-scale foundation models remains an unsolved challenge. Additionally, these methods lack the flexibility to choose the most optimal prompt at each timestep.
Understanding the latent space.
Tangential to our work, there has been substantial work [16, 1, 12] on exploring the latent space. These investigations have yielded valuable insights and practical solutions for downstream tasks such as image editing and manipulation applications [44, 34, 21, 43, 29]. Furthermore, a deeper understanding of the latent space within diffusion models has led to advancements in various methods [31, 21, 40].
3 Background
3.1 The Black Scholes Model
In this section, we provide a brief overview of the Black Scholes pricing model [26] used to determine the price of European call options of assets. which is used to determine the price of European call options on assets. In simple terms, a European call option allows an investor to lock in the price of an asset at any time but permits stock purchase (if desired) only upon expiration. Regardless of whether the stock price moves favorably or unfavorably over time, this option structure remains consistent. Investors rely on the Black-Scholes model to predict stock prices over time and make informed decisions about the optimal timing for stock purchases. The Black Scholes formula involves $5$ key variables:
  1. 1.
    Underlying stock price or spot price $S$: This represents the current price of the asset.
  2. 2.
    Strike price $K$: The strike price is the cost of the asset at the time of expiry.
  3. 3.
    Time to expiration $t$: It measures the time difference between the current moment and the expiry time.
  4. 4.
    Volatility $\sigma$: Volatility reflects the variation in prices of the asset.
  5. 5.
    Risk free rate $r$: The risk-free rate is the minimum return on an investment when the investor faces zero risks.
To obtain the Black Scholes score of purchasing an asset, the spot price $S$ is first multiplied by the standard normal probability distribution function. From this result, to obtain the final cost $C$, the strike price $K$ multiplied by the cumulative standard distribution function is subtracted. Mathematically,
$SN(d_{1})-Ke^{-rt}N(d_{2}),$ (1)
where
$d_{1}=\frac{log\frac{S}{K}+(r+\frac{\sigma^{2}}{2})(t)}{\sigma\sqrt{t}},d_{2}=d_{1}-\sigma\sqrt{t}.$ (2)
Assumptions made by the Black Scholes Model.
The Black Scholes model makes five assumptions: (i) No dividends are paid out during the life of an option, (ii) Market movements are somewhat random, (iii) There are no transaction costs in buying the asset, (iv) The volatility and risk free rate of the underlying asset are known and constant, (v) The returns of the underlying asset are normally distributed.
3.2 Relation to Diffusion Models
The main concept behind diffusion models [13] involves iteratively adding small amounts of random Gaussian noise to transform an initial photorealistic image $x_{0}$ to noise $x_{T}\sim\mathcal{N}(0,I)$ in $T$ steps. This process is known as the forward process. The converse of the theorem states that by starting from random noise $x_{T}\sim\mathcal{N}(0,I)$ and refining it iteratively for $T$ steps, we can generate a photorealistic image $x_{0}$. Since diffusion is a gradual process, the value of $T$ is typically large. At each intermediate timestep, $t\in\{0....T\}$, $x_{t}$ satisfies $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$. $0=\alpha_{T}<\alpha_{T-1}....<\alpha_{1}<\alpha_{0}=1$ are the hyperparameters of the diffusion schedule [13] and $\epsilon_{T}\sim\mathcal{N}(0,I)$. To obtain $x_{t-1}$ at each refinement step, the neural network $f_{\theta}(x_{t},t)$ is applied along with the corresponding random Gaussian noise perturbation. In other words, during each step of diffusion denoising, the known variance in added noise follows a Gaussian distribution [13]. In finance, the volatility refers to the standard deviation in the way stock prices change from one time instant to the next. The mathematical description of diffusion models, derived from non-equilibrium thermodynamics  [35], shares similarities with the derivation of the Black-Scholes model [26] used in pricing European call options within financial markets. Both models emerge from similar assumptions and conditions, underpinned by a shared mathematical structure. Assumptions (i) and (iii) made by the Black Scholes model, which talk about dividends and transaction costs, are irrelevant to diffusion models.
Figure 2: We present more results of our method along with comparisons. Vanilla SD fails to capture clear characteristics of individual text prompts, omitting distinct features such as those related to avocado/raccoon, muffin, parrot, and oil painting style. Linear interpolation generates images not consistent with the prompts, due to issues with non-linear manifolds. CLIP Min. generates images biased towards one of the prompts. Alt. and Step prompt selection methods suffer from artifacts and are not very successful in blending the characteristics of objects corresponding to the individual text prompts - the avocado/raccoon, muffin/dog are not blended well. In the parrot/dog image, the characteristics of the parrot are missing. Alt. generates artifacts in the Times Square/ oil painting image, while Times Square is not characterized well in the image generated by Step. In contrast, the Black-Scholes model adeptly overcomes these limitations, generating realistic images that meticulously balance and preserve the unique characteristics of each individual text prompt. The images are from set 1, set 2, set 3 and set 4 (Refer Section 5.1) respectively.
4 Prompt Mixing using Black Scholes Model
4.1 Problem Formulation
Consider a set of $N$ text prompts denoted as $P_{1},...P_{N}$ and $T$ diffusion denoising steps. Note that text-to-image diffusion architectures often include a text encoder that maps the text prompt to a joint text-image space. Starting from random noise and conditioned on the embedding in the text-image space, the neural network generates the final image.
Our objective is to generate an image that aligns with multiple text prompts simultaneously i.e. at the intersection of the various text-image manifolds of the diffusion models. We pose this task as a prompt mixing [27] problem. During each step of diffusion denoising, our aim is to select the most relevant text prompt (or the text prompt with respect to which the image requires further refining) for conditioning the model, ensuring the generation of an optimal image satisfying all individual text prompts.
4.2 Method
Selecting the most effective text prompt to condition a text-to-image diffusion model can be approached in various ways. Naively switching [18, 27] between prompts is sub-optimal due to the significant human effort involved. Moreover, it fails to address the critical issue of prioritizing the most relevant prompt in an automatic manner. Alternatively, one can compute the CLIP score at each step and move toward the text prompt with the lowest CLIP score. While this approach is reasonable, it overlooks essential factors related to the dynamics [13] of the diffusion denoising process, which significantly impact the final image generation. Leveraging the Black Scholes model allows us to model these dynamics effectively.
We utilize the CLIP score to quantify the “price” of the generated image, treating it as analogous to a stock. The CLIP score has been widely used to measure text-image alignment, making it a suitable metric for evaluating how well our generated image aligns with each text prompt. Consider diffusion denoising at step $i$ out of total steps $T$. We derive the relevant variables for Black Scholes formula as follows:
  1. 1.
    Underlying stock price $S$: The underlying stock price corresponds to the current value of the asset. Let $z_{t}$ represent the predicted latents at timestep $t$. We denote $z_{0,t}$ as the latents of the final predicted image extrapolated from $z_{t}$. In other words, if the denoising process were to proceed in a straightforward manner following the same direction as computed for $z_{t}$, the latents of the final predicted image would be $z_{0,t}$. The underlying stock price, denoted as $S$, is determined by the CLIP score with respect to text prompt $P_{p}$. This CLIP score can be computed using the image decoded from $z_{0,t}$. To maintain values within the range of 00 to $100$, we multiply the CLIP score by $100$.
  2. 2.
    Strike price $K$: This represents the estimated asset price at expiry, following the last step of diffusion denoising, which predicts the final image. Based on our experience with diffusion models, where we found that a CLIP score of approximately $0.25$ indicates reasonable text-image alignment, we opted for a constant value of $0.25$.
  3. 3.
    Time to expiration $t$: $t=T-i$.
  4. 4.
    Volatility $\sigma$: We compute the volatility as the square root of the variance used by the diffusion denoising scheduler at timestep $i$.
  5. 5.
    Risk-free rate $r$: $r=1/T$.
During each iteration $i$ of diffusion denoising, we calculate the Black-Scholes score $b_{i,p}$ for each text prompt $P_{p}$, $p=1...N$ using the given variables and Equation 1 and Equation 2. Subsequently, during the next step of diffusion denoising, we condition on the text prompt with the lowest Black-Scholes score. This approach ensures that the denoising process prioritizes the text prompt associated with the minimum Black-Scholes score for generating an optimal image consistent with all prompts. Please refer to Algorithm 1 for a step-wise description of the method.
Algorithm 1 Black Scholes prompt mixing in backward diffusion enables the diffusion model generate an optimal image at the intersection of multiple text-image manifolds.
1:Initialize latents to random Gaussian noise. $z_{T}\sim\mathcal{N}(0,I)$; T is the total number of diffusion denoising steps.
2:Use the text encoder $\epsilon$ to compute the CLIP embeddings of a (linguistic) combination of the text prompts $\{P\}$. $e\leftarrow\epsilon_{p}(\cup\{P\})$.
3:Initialize Black Scholes variables, strike price $K=0.25\times 100$, $rater=1/T$.
4:// Diffusion denoising - image prediction/generation loop
5:for $t\leftarrow T$ to 00 do
6:     $z_{t-1}=z_{t}-f(z_{t},t,e)$; f is the diffusion UNet.
7:     Use $z_{t-1}$ to compute $z_{0,t-1}$.
8:     Variance $\sigma$ at step $t-1$leftarrow computed using the scheduler of the diffusion model.
9:     for $i\leftarrow 1$ to $N$ do
10:         Spot price $S$leftarrow CLIP score With respect to text prompt $P_{i}$
11:         Time to expiration $\leftarrow t$
12:         Black Scholes score $b_{t,i}$ with respect to prompt $P_{i}$, at timestep $t$, $\leftarrow$ use Equation 1,2.
13:     end for
14:     $P_{min}$leftarrow Text prompt corresponding to min$\{B_{t,i}\}$, $i=1...N$
15:     $e$leftarrow $\epsilon_{p}(P_{min})$
16:end for
5 Experiments and Results
Metrics.
In line with previous research [9], we assess performance using the CLIP Score. We use two variants of the CLIP Score: (i) CLIP-combined, which computes the score by comparing the generated image against a combination of individual text prompts; and (ii) CLIP-add, which averages the CLIP scores for the generated image across each individual text prompt. To compute our metrics, we generate five images for each text prompt and use them all in the computation of our metrics.
Baselines.
Our study examines several baselines: (i) Vanilla SD: Prompt engineering, we use the vanilla stable diffusion model and condition it on a single text prompt effectively describing all individual text prompts, (ii) Lin. Interp.: Direct combination of text embeddings, achieved through linear interpolation between text embeddings [17]. This method equally weights the text embeddings associated with each text prompt. (iii) Switching between text prompts. We consider two variations here: (iii-a) Step: Following Patashnik et. al. [27], we use one text prompt for the initial 7th to 17th denoising steps and then switch to the other text prompt for the remaining steps, (iii-b) Alt.: Following Kothandaraman et. al. [18], we alternate between the two text prompts. (iv) CLIP Min.: Score-based combination, denoted as CLIP-min, where we select the text prompt corresponding to the lowest CLIP score from the previous denoising iteration.
Backbone architecture.
We use the Stable Diffusion 2.1 backbone model in all our experiments. There is no training involved, we use the pretrained model to directly perform inference.
Method CLIP-combined CLIP-add
Vanilla $0.3463$ $0.3012$
Linear Int. $0.2885$ $0.2778$
Alt. Samp. $0.3445$ $0.3098$
CLIP Min. $0.3195$ $0.2955$
Step $0.322$ $0.2997$
Black Scholes $0.3469$ $0.3112$
Table 1: We follow prior work [9, 3] and analyze using CLIP Scores. The Black Scholes algorithm for prompt mixing in diffusion models achieves superior results, compared to other prompt-mixing techniques, as also evidenced by the qualitative results.
5.1 Experimental Settings
We consider the following experimental settings:
Set 1:
Simple text prompts with single objects. In this experiment, we validate our approach using straightforward scenarios, we use two text prompts, each describing one class. Specifically, we consider $17$ classes: [‘a rock’, ‘a coffee mug’, ‘a cute dog’, ‘a pink cat’, ‘a teddy bear’, ‘a robot’, ‘an alien’, ‘an avocado’, ‘a cute raccoon’, ‘a corgi dog’, ‘a parrot’, ‘a car’, ‘a squirrel’, ‘a cute rabbit’, ‘pizza’, ‘muffin’, ‘icecream’], and construct $13617c_{2}=$ 136 text prompts.
Set 2:
Multiple objects per text prompt. To add complexity, blend multiple objects, in the presence of additional objects in the scene. The goal of this experiment is to assess the models’ capabilities in blending the right objects in the scene, and preserving the characteristics of the other objects. We work with three classes of objects: [‘a basket’, ‘a teapot’], [‘apples’, ‘bananas’, ‘a cute cat’, ‘a cute dog’, ‘muffins’], [‘a table’, ‘a carpet’, ‘a bed’] and construct combinations using the first and second set and the second and third set. Additionally, we include the data point “a cat in a bathtub” and “a corgi dog in a bathtub,” resulting in a total of $71$ prompts.
Set 3:
Object actions against backgrounds (prompt mixing w.r.t. object). We investigate the model’s ability to morph objects while considering object-level action information and scene or background context. We consider the following backgrounds, [‘walking in Times Square’, ‘skiing in Times Square’, ‘walking in a beautiful garden’, ‘surfing on the beach’, ‘eating watermelon on the beach’, ‘sitting on a sofa on the beach’, ‘watching northern lights’, ‘watching sunset at a beach’, ‘admiring the opera house in Sydney’, ‘sleeping in a cozy bedroom’, ‘admiring a beautiful waterfall in a forest’, ‘walking in a cherry blossom garden’, ‘walking in a colorful autumn forest’, ‘flying in the sky at sunset’]. The objects are [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. In total, we create $210$ prompts.
Set 4:
Object performing action against a background (prompt mixing w.r.t. background). In this experiment, we explore how the model blends global information and backgrounds while considering objects in various scenarios. We consider the objects [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. The backgrounds and style prompts are [‘walking in Times Square’, ‘sunset’, ‘van gogh style’, ‘oil painting’, ‘a garden with tulips’, ‘northern lights’]. Overall, we have $90$ prompts in this experiment.
5.2 Analysis and Comparisons
We show qualitative results in Figure 1 and Figure 2. More qualitative results can be found in the appendix. The quantitative comparisons are in Figure 1. Our quantitative results are averaged across the metrics for each of the four sets, with $680$, $355$, $1050$, $450$ images respectively. A detailed analysis of our method and benefits over prior art is as follows:
  • Vanilla stable diffusion (SD): Vanilla stable diffusion v2.1 can create plausible combinations of the provided text prompts. However, due to hallucination issues, it is constrained by the distributions it learned during training. As a result, it struggles to capture clear characteristics of individual text prompts. For example, in Figure 1, the distinct features of the parrot, cat, dog/penguin, and sunset/penguin are missing. Similarly, in Figure 2, the avocado/raccoon, muffin, parrot, and oil painting style lack distinct characteristics. Despite this, the generated images align with the broad text prompt descriptions, resulting in a considerable CLIP score.
  • Linear interpolation [17]: Linear interpolation is ill-suited for generating an image that combines two text prompts. This limitation arises from the non-linear nature of the text-image space. While linear interpolation assumes a simple linear relationship, the actual mapping between textual descriptions and image features is intricate. Furthermore, linear interpolation can only traverse a straight path between two points in the latent space, failing to capture nuanced variations or produce novel features beyond the endpoints.
  • Alternating Sampling [18],: Alternating sampling generates features related to both text prompts in the final output image. However, the objects produced suffer from poor definition, unrealistic appearance, and low quality in many instances. Artifacts are prevalent, and several generated images appear implausible. The underlying issue lies in the algorithm’s routine alternation between the two text prompts during diffusion denoising, without adequately emphasizing the most relevant prompt at each step. Despite these limitations, the model achieves a reasonable CLIP Score by broadly aligning with the overall image description, akin to vanilla denoising.
  • Step-wise switching [18]: Step-wise switching is similarly ineffective for the same reasons as Alternating sampling.
  • CLIP-min: The CLIP-min results exhibit bias toward one of the text prompts in most cases, preventing the model from generating an image that aligns with the intersection of both prompts. This bias arises because the model overlooks the dynamics of diffusion denoising and instead selects the text prompt with the lowest CLIP score at each step, resulting in bias-related issues.
  • Black Scholes algorithm: Our approach based on the Black-Scholes algorithm effectively generates realistic images that align with the intersection of the two text prompts. These images exhibit minimal unrealistic artifacts. Additionally, our algorithm successfully preserves individual characteristics corresponding to each text prompt. By modeling the dynamics of diffusion denoising, the algorithm strategically selects the optimal text prompt at each step, achieving a balanced synthesis of both prompts. Notably, our model attains superior quantitative results.
In summary, the Black-Scholes model outperforms all previous methods for prompt mixing by generating realistic images that meticulously preserve and balance the characteristics of each individual text prompt.
Figure 3: Image variations from our method using the Black Scholes model for a single prompt, starting from different random Gaussian noise initializations.
6 Conclusions, Limitations, and Future Work
In this paper, we introduce a novel perspective on prompt mixing, inspired by a probablistic prediction model and its theoretical framework from the domain of Finance. We draw parallels between the Black-Scholes algorithm used in financial markets and diffusion models. From this comparison, we derive an algorithm for prompt mixing, aiming to generate images that align with multiple text prompts. Our extensive analysis showcase both qualitative and quantitative advantages of our approach. Our study has a few limitations, which opens avenues for future work: (i) We follow prior work [9, 3] by using CLIP scores for quantitative evaluation. However, CLIP scores may not be the optimal evaluation strategy, as they can yield similar scores for images with significant quality differences. Due to its learned nature, CLIP is vulnerable to hallucination and bias issues related to its training data. Consequently, it may assign similar scores to images that exhibit general characteristics of the text prompt, without focusing on specific details and features of the scene entities described in the text prompt. To address this, future research could explore more advanced image-language models [25] for evaluation; the development of good metrics for the evaluation of generative models is a widespread issue, and could help better quantify the performance of our method. (ii) Our study focuses on prompt mixing with two prompts. Investigating the impact of using more than two prompts could be a promising direction for future work. (iii) While our method leverages diffusion models, it may not be applicable to non-Gaussian [6] diffusion models or one-step diffusion models [41], as demonstrated in recent research.
Moreover, future work on integrating our approach with recent advancements in attention guidance, layout modeling, and related areas could extend its applicability to critical downstream tasks such as image editing [17, 39, 4], compositionality [24, 3], handling complex prompts [23], text-based view synthesis [18, 19] and personalized image generation [32, 20]. While recognizing the success of recent text-to-image models like Dalle-3, Stable Diffusion 3, Firefly 3, and Imagen 3 in generating images from diverse and complex text prompts, we introduce a data-efficient approach to prompt mixing. Our method offers advantages over previous techniques for prompt-mixing, and future extension could explore integrating our model’s benefits into state-of-the-art text-to-image architectures for data efficient prompt mixing and downstream applications.
Societal impact.
While our method offers valuable tools for generative AI-based content creation, its potential misuse underscores the need for research in watermarking and deepfake detection to mitigate risks.
Acknowledgements.
This work was supported in part by ARO Grants W911NF2110026, W911NF2310046, W911NF2310352 and Army Cooperative Agreement W911NF2120076.

### v2

@@
-Prompt Mixing in Diffusion Models using the Black Scholes Algorithm
+Financial Models in Generative Art: Black-Scholes-Inspired Concept Blending in Text-to-Image Diffusion
 Abstract
-We introduce a novel approach for prompt mixing, aiming to generate images at the intersection of multiple text prompts using pre-trained text-to-image diffusion models.
+We introduce a novel approach for concept blending in pretrained text-to-image diffusion models, aiming to generate images at the intersection of multiple text prompts.
 At each time step during diffusion denoising, our algorithm forecasts predictions w.r.t.
 the generated image and makes informed text conditioning decisions.
-To do so, we leverage the connection between diffusion models (rooted in non-equilibrium thermodynamics) and the Black-Scholes model for pricing options in Finance, and draw analogies between the variables in both contexts to derive an appropriate algorithm for prompt mixing using the Black Scholes model.
-Specifically, the parallels between diffusion models and the Black-Scholes model enable us to leverage properties related to the dynamics of the Markovian model derived in the Black-Scholes algorithm.
-Our prompt-mixing algorithm is data-efficient, meaning it does not need additional training.
+Central to our method is the unique analogy between diffusion models, which are rooted in non-equilibrium thermodynamics, and the Black-Scholes model for financial option pricing.
+By drawing parallels between key variables in both domains, we derive a robust algorithm for concept blending that capitalizes on the Markovian dynamics of the Black-Scholes framework.
+Our text-based concept blending algorithm is data-efficient, meaning it does not need additional training.
 Furthermore, it operates without human intervention or hyperparameter tuning.
-We highlight the benefits of our approach by comparing it qualitatively and quantitatively to other prompt mixing techniques, including linear interpolation, alternating prompts, step-wise prompt switching, and CLIP-guided prompt selection across various scenarios such as single object per text prompt, multiple objects per text prompt and objects against backgrounds.
+We highlight the benefits of our approach by comparing it qualitatively and quantitatively to other text based concept blending techniques, including linear interpolation, alternating prompts, step-wise prompt switching, and CLIP-guided prompt selection across various scenarios such as single object per text prompt, multiple objects per text prompt and objects against backgrounds.
+Our work shows that financially inspired techniques can enhance text-to-image concept blending in generative AI, paving the way for broader innovation.
 Code is available at https://github.com/divyakraman/BlackScholesDiffusion2024.
@@
 1 Introduction
-Text-to-image diffusion models [13, 14] such as Imagen [33] by Google, Firefly [2] by Adobe, DALLE [7] by Microsoft, and Stable Diffusion [31, 11] by Stability AI can produce photorealistic images from textual descriptions and are typically trained on vast amounts of data using substantial computational resources.
-The latest versions of these models (as of May 2024) have seen significant improvements in comparison to their initial versions, attributed to various factors, such as increased training data, enhanced text understanding networks, improved image generation networks, and optimized strategies.
-Data efficient solutions [36], leveraging pre-trained diffusion models, can enhance results without requiring additional data collection.
-Prompt engineering [37] is one effective approach, involving techniques such as rephrasing prompts to improve model generalization.
-Another method is using large language models (LLMs) [23, 38] to parse complex prompts and determine useful priors for image generation.
-Attention maps [9] and prompt mixing techniques [27, 5] also contribute to better results.
-Prompt mixing becomes particularly valuable when aiming to generate images that blend different text-image concepts i.e.
-when aiming to generate images at the intersection [27, 30, 28] of different text-image concepts.
-For instance, consider generating an image resembling a hybrid of a pink cat and a dog.
-Current diffusion models [2, 33, 7, 11] excel at straightforwardly combining individual text prompts to create visually appealing combinations.
-However, these capabilities are often restricted to advanced model versions (version 3) and rely on large-data techniques.
-Data efficient prompt mixing techniques can enhance image generation without additional data.
-While linear interpolation [17] of text embeddings from individual prompts is a simple approach, it may not be optimal due to the highly non-linear nature of the text-image manifold and potential bias issues.
-Alternatives include switching between prompts during diffusion denoising, either by alternating [18] or using a step-wise technique [27].
-However, these methods often require human involvement and careful prompt engineering for optimal results.
-The question arises: what’s the best way to switch between text prompts for prompt-mixing?
-An automated approach considers the model’s varying capabilities with different prompts.
-While attention maps [15, 9] and layout guidance![42, 10] methods excel at guiding the model toward distinct scene entities, they may not be as effective for blending concepts within the same entity.
-To address this, during each step of diffusion denoising, the network should automatically focus on the text prompt that corresponds to the image’s deficiencies.
+Text-to-image diffusion models [13, 14] have become essential tools for creative applications such as content creation, editing, personalization, and zero-shot generation.
+However, a core challenge remains: generating new and unseen concepts without retraining the model or collecting additional task-specific data.
+One promising avenue is to leverage the model’s intrinsic understanding of various conceptual manifolds to synthesize novel outputs by operating at their intersection.
+This process, known as concept blending or prompt mixing [26, 5, 29, 27], used interchangeably in this paper, involves integrating two distinct concepts to produce a single, coherent image.
+For example, one could generate an image that merges the features of a pink cat with those of a dog.
+Although state-of-the-art diffusion models [2, 32, 7, 11] can blend text prompts into visually appealing results, these capabilities typically depend on advanced architectures and large-scale training data.
+In contrast, data-efficient prompt mixing techniques hold the potential to enhance image generation while avoiding extra data collection costs.
+While simple strategies like linear interpolation of text embeddings [17] exist, they may not be optimal given the highly nonlinear nature of the text-image manifold and the risks of bias.
+Other methods, such as alternating [18] or step-wise prompt switching [26], encounter challenges stemming from human involvement and meticulous prompt engineering.
+The central question we address is: How can we optimally switch between text prompts for effective concept blending?
+An ideal solution would automatically account for the varying capabilities of the model with respect to different prompts.
+While attention maps [15, 9] and layout guidance [40, 10] excel at directing the model toward distinct scene entities, they are less effective for blending concepts within the same entity.
+To overcome this, we propose an automated mechanism in which, at each step of the diffusion denoising process, the model focuses on the text prompt most aligned with the current image deficiencies.
 Main Contributions.
-In this paper, we introduce a novel approach to prompt mixing by integrating concepts from economics and finance.
-Our objective is to leverage pretrained diffusion models for prompt-mixing to generate images at the intersection of various text-image manifolds.
-During each step of the diffusion denoising process, our method dynamically conditions on the text prompt that requires the highest level of attention.
-To achieve this, our algorithm assesses the ‘cost’ associated with each relevant text prompt and selects the optimal conditions for the subsequent step of image generation based on the ‘cost’ that requires maximum optimization.
-Key Insight: The inspiration for our approach lies in non-equilibrium thermodynamics [35], which is at the core of the development of diffusion models.
-These models share a conceptual foundation with the Black Scholes model[26], a Nobel Prize-winning mathematical framework extensively employed in financial markets for pricing European call options[8].
-The denoising steps within a diffusion model, aimed at image generation, constitute a Markovian time-series.
-In our analogy, the image itself represents a valuable ‘stock’ or ‘asset’ that we aim to ‘purchase’ (or generate) at the most favorable ‘cost’ (or alignment with the text prompts).
-During each step of the diffusion denoising process, we extrapolate the generate latents to compute the ‘stock’ prices at the corresponding time-step, and use the properties of diffusion models to compute the various variables involved in the Black Scholes algorithm.
-Consequently, we leverage the Black Scholes model to predict a score for each text prompt.
-This score serves as an indicator of how the image should be conditioned in the subsequent timestep.
-By adopting this approach, our model dynamically focuses on aspects that require attention, ultimately generating an image that optimally satisfies all relevant text prompts.
-We perform experiments on prompts with varying complexities to assess our method’s ability to seamlessly blend different objects and backgrounds across scenarios.
-Through qualitative and quantitative comparisons against several baselines-such as linear interpolation, prompt switching, and CLIP score-we highlight the superiority of our approach.
+In this paper, we introduce a novel approach for concept blending that incorporates ideas from economics and finance.
+Our method leverages pre-trained diffusion models to generate images that reside at the intersection of various text-image manifolds, dynamically conditioning the generation process on the prompt demanding the greatest attention.
+Specifically, our algorithm evaluates the “cost” associated with each text prompt and selects the optimal conditioning for the subsequent generation step based on which prompt requires the most optimization effort.
+Key Insight: Our approach is inspired by non-equilibrium thermodynamics [34]—the foundational principle behind diffusion models—and shares a conceptual basis with the Black–Scholes model [25, 8], a celebrated framework in financial markets for pricing European stock options.
+In our analogy, the image functions as a valuable asset whose generation should occur at the most favorable “cost,” determined by its alignment with the input text prompts.
+Concretely, at each denoising step we:
+  1. 1.
+    Extrapolate the latent representations to compute analogous “stock prices” at each timestep.
+  2. 2.
+    Leverage structural properties of diffusion models to derive the variables in the Black–Scholes algorithm.
+  3. 3.
+    Predict a score for each text prompt, indicating how the image should be conditioned in subsequent steps.
+This mechanism enables our model to adaptively concentrate on the aspects most in need of refinement, ultimately generating images that optimally reflect all relevant text inputs.
+We validate our approach through extensive experiments on prompts of varying complexity, demonstrating the model’s ability to seamlessly blend different objects and backgrounds.
+Quantitative and qualitative comparisons against baselines such as vanilla Stable Diffusion, linear interpolation, alternating sampling, prompt switching, and CLIP-guided prompt switching confirm the superiority of our method over current state-of-the-art techniques.
+Our work demonstrates that integrating financial methodologies into text-to-image synthesis can significantly improve concept blending, laying the groundwork for broader innovations in generative AI.
 2 Related work
-Prompt mixing.
-Prompt mixing is a technique where different text prompts are used at various steps during the diffusion denoising process.
-Patashnik et al.
-[27] follow a step-wise approach, while Aerial Diffusion [18] alternates between prompts to create semantically consistent aerial-view images.
-Tools such as Image Mixer Diffusion [28] and CLIP Guided Image Mixing [30] follow similar techniques of prompt switching to blend multiple text prompts.
-However, these methods often require complex hyperparameter tuning.
-Mixing step.
-The mixing time or mixing step [22] of a Markov chain refers to the duration it takes for the chain to reach its steady-state distribution.
-Mixing time has found applications in image editing [46] and synthesizing out-of-distribution (OOD) images [45].
-When switching between prompts using mixing time, a step-wise approach is followed, akin to the work by Patashnik et al.
-[27].
-The advantage of mixing-time approaches over the method proposed by Patashnik et al.
-[27] is that it mathematically determines the optimal switching time without requiring complex hyperparameter tuning.
-To approximate the mixing time, it estimates the radius of the latent space.
-However, there are a few limitations to these methods.
-Firstly, most of these approaches are applied to diffusion models trained on relatively small, problem-specific datasets.
-Secondly, identifying clear boundaries on large-scale foundation models remains an unsolved challenge.
-Additionally, these methods lack the flexibility to choose the most optimal prompt at each timestep.
-Understanding the latent space.
-Tangential to our work, there has been substantial work [16, 1, 12] on exploring the latent space.
-These investigations have yielded valuable insights and practical solutions for downstream tasks such as image editing and manipulation applications [44, 34, 21, 43, 29].
-Furthermore, a deeper understanding of the latent space within diffusion models has led to advancements in various methods [31, 21, 40].
-3 Background
+2.1 Concept Blending / Prompt Mixing
+Prompt engineering [35] employs techniques such as rephrasing prompts to enhance model generalization.
+Additionally, large language models (LLMs) [23, 36] have been used to parse complex prompts and extract useful priors for image generation.
+Complementary to these methods are techniques involving attention maps [9] and prompt mixing, which applies different text prompts at various steps during the diffusion denoising process.
+For instance, Patashnik et al.
+[26] adopt a step-wise approach, whereas Aerial Diffusion [18] alternates prompts to produce semantically consistent aerial-view images.
+Similar strategies are employed by tools such as Image Mixer Diffusion [27] and CLIP Guided Image Mixing [29].
+A common drawback is that these methods often necessitate complex hyperparameter tuning.
+2.2 Mixing Step
+The mixing time (or mixing step) of a Markov chain, as discussed in [22], represents the time required for the chain to reach its steady-state distribution.
+This concept has been exploited in areas such as image editing [44] and the synthesis of out-of-distribution (OOD) images [43].
+When using mixing time for prompt switching—similar to the step-wise methods of Patashnik et al.
+[26]—one key advantage is that the optimal switching time is determined mathematically, thereby reducing the need for extensive hyperparameter tuning.
+Typically, the mixing time is approximated by estimating the radius of the latent space.
+However, these approaches often apply to diffusion models trained on relatively small, problem-specific datasets and may struggle with the challenges of larger-scale foundation models.
+Moreover, they generally lack the flexibility to select the most optimal prompt at every timestep.
+2.3 Understanding the Latent Space
+Exploration of the latent space has been a significant area of investigation [16, 1, 12].
+Insights from these studies have facilitated advancements in downstream tasks such as image editing and manipulation [42, 33, 21, 41, 28].
+Furthermore, a deeper understanding of the latent space in diffusion models has spurred improvements in various methods [30, 21, 38], providing a foundation for more effective image synthesis and manipulation techniques.
+3 The Black Scholes Algorithm and Diffusion Models
 3.1 The Black Scholes Model
 In this section, we provide a brief overview of the Black Scholes pricing model [25] used to determine the price of European call options of assets.
-which is used to determine the price of European call options on assets.
 In simple terms, a European call option allows an investor to lock in the price of an asset at any time but permits stock purchase (if desired) only upon expiration.
@@
 $d_{1}=\frac{log\frac{S}{K}+(r+\frac{\sigma^{2}}{2})(t)}{\sigma\sqrt{t}},d_{2}=d_{1}-\sigma\sqrt{t}.$ (2)
-Assumptions made by the Black Scholes Model.
-The Black Scholes model makes five assumptions: (i) No dividends are paid out during the life of an option, (ii) Market movements are somewhat random, (iii) There are no transaction costs in buying the asset, (iv) The volatility and risk free rate of the underlying asset are known and constant, (v) The returns of the underlying asset are normally distributed.
 3.2 Relation to Diffusion Models
-The main concept behind diffusion models [13] involves iteratively adding small amounts of random Gaussian noise to transform an initial photorealistic image $x_{0}$ to noise $x_{T}\sim\mathcal{N}(0,I)$ in $T$ steps.
+3.2.1 Diffusion Models
+The main concept behind diffusion models involves iteratively adding small amounts of random Gaussian noise to transform an initial photorealistic image $x_{0}$ into noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ steps.
 This process is known as the forward process.
-The converse of the theorem states that by starting from random noise $x_{T}\sim\mathcal{N}(0,I)$ and refining it iteratively for $T$ steps, we can generate a photorealistic image $x_{0}$.
-Since diffusion is a gradual process, the value of $T$ is typically large.
-At each intermediate timestep, $t\in\{0....T\}$, $x_{t}$ satisfies $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$.
-$0=\alpha_{T}<\alpha_{T-1}....<\alpha_{1}<\alpha_{0}=1$ are the hyperparameters of the diffusion schedule [13] and $\epsilon_{T}\sim\mathcal{N}(0,I)$.
+Conversely, starting from random noise $x_{T}\sim\mathcal{N}(0,I)$ and refining it iteratively for $T$ steps can generate a photorealistic image $x_{0}$.
+Since diffusion is gradual, $T$ is typically large.
+At each intermediate timestep $t\in\{0,\ldots,T\}$, $x_{t}$ satisfies:
+$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$
+The hyperparameters of the diffusion schedule are $10=\alpha_{T}<\alpha_{T-1}<\ldots<\alpha_{1}<\alpha_{0}=$ 1, and $\epsilon_{T}\sim\mathcal{N}(0,I)$.
 To obtain $x_{t-1}$ at each refinement step, the neural network $f_{\theta}(x_{t},t)$ is applied along with the corresponding random Gaussian noise perturbation.
-In other words, during each step of diffusion denoising, the known variance in added noise follows a Gaussian distribution [13].
-In finance, the volatility refers to the standard deviation in the way stock prices change from one time instant to the next.
-The mathematical description of diffusion models, derived from non-equilibrium thermodynamics  [35], shares similarities with the derivation of the Black-Scholes model [26] used in pricing European call options within financial markets.
-Both models emerge from similar assumptions and conditions, underpinned by a shared mathematical structure.
-Assumptions (i) and (iii) made by the Black Scholes model, which talk about dividends and transaction costs, are irrelevant to diffusion models.
+Essentially, during each step of diffusion denoising, the known variance in added noise follows a Gaussian distribution.
+3.2.2 Analysis based on Thermodynamics
+The mathematical formulation of diffusion models, which is rooted in non-equilibrium thermodynamics, exhibits striking similarities to the derivation of the Black-Scholes model used for pricing European call options in financial markets.
+Both models arise from analogous assumptions and share a common mathematical structure.
+Diffusion Models
+In the context of diffusion, tools from statistical mechanics reveal that the generative dynamics undergo phase transitions and symmetry breaking.
+The dynamic equation of the generative process can be interpreted as a stochastic adiabatic transformation that minimizes free energy while maintaining thermal equilibrium.
+Moreover, diffusion models incorporate Gaussian noise in a controlled manner to learn the underlying data distribution: forward process, driving the system out of equilibrium by adding noise and reverse process, guiding the system back to a structured state.
+The Black-Scholes Model
+The Black-Scholes model assumes that the stock price $S_{t}$ follows a geometric Brownian motion, described by a stochastic differential equation (SDE).
+With the application of Itô’s Lemma, one derives a partial differential equation (PDE) that characterizes the evolution of option prices.
+Here, the pricing of derivatives can be viewed as a process of minimizing a form of “financial free energy” under specific constraints, analogous to thermodynamic systems seeking equilibrium.
+A Unified Thermodynamic View
+Both models can be interpreted through the lens of thermodynamics:
+  • In diffusion models, the reverse process minimizes a free energy functional:
+[…]
+    where $q(\mathbf{x})$ represents the target distribution.
+  • In the Black-Scholes framework, option pricing is analogous to minimizing a financial free energy, where the equilibrium state corresponds to a fair, risk-neutral price.
+Thus, while diffusion models and the Black-Scholes equation address different applications—one in generative modeling and the other in finance—they both embody the principles of free energy minimization and equilibrium.
+The forward process in diffusion models perturbs the system away from equilibrium using controlled noise, whereas the Black-Scholes approach interprets market fluctuations through Brownian motion, seeking an equilibrium price despite the inherent randomness in financial markets.
+3.2.3 Analysis based on SDEs
+Both diffusion models and the Black-Scholes equation rely on stochastic differential equations (SDEs) to describe the evolution of systems over time.
+This common mathematical framework allows us to draw interesting parallels between seemingly disparate domains like image generation and financial option pricing.
+Diffusion Models
+In text-to-image diffusion models, the forward diffusion process is governed by the SDE:
+$\mathbf{x}_{t}=-\frac{1}{2}\beta_{t}\mathbf{x}_{t}\,dt+\sqrt{\beta_{t}}\,d\mathbf{w}_{t},$
+where $\mathbf{x}_{t}$ is the state (i.e., the image representation) at time $t$, $\beta_{t}$ is a time-dependent noise coefficient, and $d\mathbf{w}_{t}$ denotes the increment of a Wiener process (Brownian motion).
+To reconstruct the data, the model uses the reverse diffusion process:
+$\mathbf{x}_{t}=\left(\frac{1}{2}\beta_{t}\mathbf{x}_{t}+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right)dt+\sqrt{\beta_{t}}\,d\mathbf{w}_{t},$
+where $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is the score function, guiding the reverse process.
+Black-Scholes Equation
+In finance, the Black-Scholes framework models the stock price $S_{t}$ as following a geometric Brownian motion:
+$dS_{t}=\mu S_{t}\,dt+\sigma S_{t}\,dW_{t},$
+with $\mu$ as the drift rate, $\sigma$ as the volatility, and $dW_{t}$ as the Wiener process.
+Using Itô’s Lemma, the evolution of the option price $V(S,t)$ is given by:
+$dV=\left(\frac{\partial V}{\partial t}+\mu S\frac{\partial V}{\partial S}+\frac{1}{2}\sigma^{2}S^{2}\frac{\partial^{2}V}{\partial S^{2}}\right)dt+\sigma S\frac{\partial V}{\partial S}\,dW_{t}.$
+This derivation leads to the well-known Black-Scholes partial differential equation (PDE):
+$\frac{\partial V}{\partial t}+rS\frac{\partial V}{\partial S}+\frac{1}{2}\sigma^{2}S^{2}\frac{\partial^{2}V}{\partial S^{2}}-rV=0,$
+where $r$ is the risk-free interest rate.
+Connecting the Two Domains
+Both approaches yield PDEs describing the evolution of key quantities:
+  1. 1.
+    In diffusion models, the Fokker-Planck equation governs the probability density $p(\mathbf{x},t)$ of the image states:
+    $\frac{\partial p(\mathbf{x},t)}{\partial t}=-\nabla\cdot\left(\mathbf{f}(\mathbf{x},t)p(\mathbf{x},t)\right)+\frac{1}{2}\nabla\cdot\left(D(\mathbf{x},t)\nabla p(\mathbf{x},t)\right),$
+    where $\mathbf{f}(\mathbf{x},t)$ is the drift term and $D(\mathbf{x},t)$ the diffusion coefficient.
+  2. 2.
+    In the Black-Scholes model, the option price $V(S,t)$ evolves according to a PDE derived from the stock’s stochastic dynamics.
+Moreover, both models use gradients to guide their respective processes.
+In diffusion models, the score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ directs the reverse diffusion process, while in the Black-Scholes framework, the gradient of $V(S,t)$ with respect to $S$ (often referred to as delta) is central to constructing a risk-neutral portfolio.
+This similarity highlights a fascinating cross-domain synergy where concepts from financial mathematics can inspire novel techniques in generative modeling, and vice versa.
 Figure 2: We present more results of our method along with comparisons.
@@
 The images are from set 1, set 2, set 3 and set 4 (Refer Section 5.1) respectively.
-4 Prompt Mixing using Black Scholes Model
+4 Blending Concepts using the Black Scholes Algorithm
 4.1 Problem Formulation
@@
 at the intersection of the various text-image manifolds of the diffusion models.
-We pose this task as a prompt mixing [27] problem.
+We pose this task as a prompt mixing or concept blending [26] problem.
 During each step of diffusion denoising, our aim is to select the most relevant text prompt (or the text prompt with respect to which the image requires further refining) for conditioning the model, ensuring the generation of an optimal image satisfying all individual text prompts.
@@
 Selecting the most effective text prompt to condition a text-to-image diffusion model can be approached in various ways.
-Naively switching [18, 27] between prompts is sub-optimal due to the significant human effort involved.
-Moreover, it fails to address the critical issue of prioritizing the most relevant prompt in an automatic manner.
-Alternatively, one can compute the CLIP score at each step and move toward the text prompt with the lowest CLIP score.
-While this approach is reasonable, it overlooks essential factors related to the dynamics [13] of the diffusion denoising process, which significantly impact the final image generation.
-Leveraging the Black Scholes model allows us to model these dynamics effectively.
-We utilize the CLIP score to quantify the “price” of the generated image, treating it as analogous to a stock.
-The CLIP score has been widely used to measure text-image alignment, making it a suitable metric for evaluating how well our generated image aligns with each text prompt.
-Consider diffusion denoising at step $i$ out of total steps $T$.
-We derive the relevant variables for Black Scholes formula as follows:
+Naively switching [18, 26] between prompts is sub-optimal—not only does it require significant human intervention, but it also fails to automatically prioritize the most relevant prompt.
+An alternative approach computes the CLIP score at each step and moves toward the prompt with the lowest CLIP score; however, this method overlooks the dynamic aspects inherent in the diffusion denoising process [13].
+To address these issues, we leverage the Black-Scholes model to effectively capture these dynamics.
+4.2.1 Black-Scholes Analogy in Diffusion Models
+Our approach uses the CLIP score to quantify the “price” of the generated image, treating it as the underlying asset in a stock pricing analogy.
+The CLIP score—which is widely used to assess text-image alignment—provides a natural metric for how well the image aligns with each text prompt.
+For diffusion denoising at step $i$ out of a total of $T$ steps, we derive the Black-Scholes variables as follows:
   1. 1.
-    Underlying stock price $S$: The underlying stock price corresponds to the current value of the asset.
-Let $z_{t}$ represent the predicted latents at timestep $t$.
-We denote $z_{0,t}$ as the latents of the final predicted image extrapolated from $z_{t}$.
-In other words, if the denoising process were to proceed in a straightforward manner following the same direction as computed for $z_{t}$, the latents of the final predicted image would be $z_{0,t}$.
-The underlying stock price, denoted as $S$, is determined by the CLIP score with respect to text prompt $P_{p}$.
-This CLIP score can be computed using the image decoded from $z_{0,t}$.
-To maintain values within the range of 00 to $100$, we multiply the CLIP score by $100$.
+    Underlying Stock Price, $S$: Analogous to the current value of an asset, we define $S$ as the CLIP score measuring the alignment of the generated image with a specific text prompt $P_{p}$ at the current stage.
+Let $z_{t}$ denote the predicted latents at timestep $t$, and $z_{0,t}$ represent the latents of the final predicted image extrapolated from $z_{t}$.
+The score is scaled (multiplied by 100) to lie within the range 00 to $100$.
   2. 2.
-    Strike price $K$: This represents the estimated asset price at expiry, following the last step of diffusion denoising, which predicts the final image.
-Based on our experience with diffusion models, where we found that a CLIP score of approximately $0.25$ indicates reasonable text-image alignment, we opted for a constant value of $0.25$.
+    Strike Price, $K$: This represents the asset’s price at expiration.
+In our formulation, $K$ is defined as the average CLIP score observed at the final diffusion step when the model generates an image using a straightforward combination of prompts.
+This value reflects the model’s maximum potential for that set of prompts.
   3. 3.
-    Time to expiration $t$: $t=T-i$.
+    Time to Expiration, $t$: In diffusion models, $t$ corresponds to the number of remaining denoising steps, i.e., $t=T-i$, where $i$ is the current step.
   4. 4.
-    Volatility $\sigma$: We compute the volatility as the square root of the variance used by the diffusion denoising scheduler at timestep $i$.
+    Volatility, $\sigma$: In finance, volatility is the standard deviation of the asset price fluctuations.
+Analogously, we compute $\sigma$ as the square root of the variance used by the diffusion denoising scheduler at timestep $i$, reflecting the uncertainty inherent in the diffusion process.
   5. 5.
-    Risk-free rate $r$: $r=1/T$.
-During each iteration $i$ of diffusion denoising, we calculate the Black-Scholes score $b_{i,p}$ for each text prompt $P_{p}$, $p=1...N$ using the given variables and Equation 1 and Equation 2.
-Subsequently, during the next step of diffusion denoising, we condition on the text prompt with the lowest Black-Scholes score.
-This approach ensures that the denoising process prioritizes the text prompt associated with the minimum Black-Scholes score for generating an optimal image consistent with all prompts.
-Please refer to Algorithm 1 for a step-wise description of the method.
-Algorithm 1 Black Scholes prompt mixing in backward diffusion enables the diffusion model generate an optimal image at the intersection of multiple text-image manifolds.
+    Risk-free Rate, $r$: In traditional finance, the risk-free rate is determined by economic factors.
+For our purposes, we simplify by setting $r=\frac{1}{T}$ to distribute returns equally over the entire diffusion process.
+4.2.2 Black-Scholes Guided Prompt Selection
+At each iteration $i$ of the diffusion denoising process, we compute the Black-Scholes score $b_{i,p}$ for each text prompt $P_{p}$ (with $p=1,\dots,N$) using the variables introduced above along with Equations 1 and 2.
+In the subsequent diffusion step, the model conditions on the text prompt associated with the lowest Black-Scholes score.
+This ensures that the denoising process prioritizes the prompt that most needs refinement, enhancing the overall alignment of the generated image with all inputs.
+For a detailed step-by-step description of our method, please refer to Algorithm 1.
+Algorithm 1 Black Scholes concept blending in backward diffusion enables the diffusion model generate an optimal image at the intersection of multiple text-image manifolds.
 1:Initialize latents to random Gaussian noise. $z_{T}\sim\mathcal{N}(0,I)$; T is the total number of diffusion denoising steps.
@@
 Metrics.
-In line with previous research [9], we assess performance using the CLIP Score.
-We use two variants of the CLIP Score: (i) CLIP-combined, which computes the score by comparing the generated image against a combination of individual text prompts; and (ii) CLIP-add, which averages the CLIP scores for the generated image across each individual text prompt.
-To compute our metrics, we generate five images for each text prompt and use them all in the computation of our metrics.
+We evaluate performance using the following metrics:
+  1. 1.
+    Text-image alignment: We utilize two variants of the CLIP score:
+    • CLIP-combined: This variant assesses the overall alignment with text by comparing the generated image against a combination of individual text prompts.
+    • CLIP-add: This averages the CLIP scores for the generated image across each individual text prompt, reflecting alignment with each specific concept.
+  2. 2.
+    Preserving fidelity (or characteristics) w.r.t. each constituent concept:
+    • The BLIP score is calculated by comparing the generated image to a combination of individual text prompts, measuring overall text-level alignment.
+    • The DINO score evaluates the generated image against each individual text prompt, assessing how well the image preserves the characteristics of each concept.
+This indicates the fidelity of the attributes in the generated image relative to the individual concepts.
+    Thus, while BLIP focuses on overall concept blending, DINO provides insights into the preservation of characteristics for each concept.
+We aim for high values in both scores, calculating a net score by multiplying the DINO and BLIP scores for each prompt and averaging across all prompts.
+  3. 3.
+    Generation quality: We use KID to assess the realism of the generated samples, serving as an indicator of their overall quality.
+To calculate our metrics, we generate five images for each text prompt and incorporate all of them into the metric computations.
 Baselines.
-Our study examines several baselines: (i) Vanilla SD: Prompt engineering, we use the vanilla stable diffusion model and condition it on a single text prompt effectively describing all individual text prompts, (ii) Lin.
-Interp.: Direct combination of text embeddings, achieved through linear interpolation between text embeddings [17].
+Our study examines several baselines: (i) Vanilla SD: Prompt engineering, we use the vanilla stable diffusion model and condition it on a single text prompt effectively describing all individual text prompts, (ii) Linear Interpolation: Direct combination of text embeddings, achieved through linear interpolation between text embeddings [17].
 This method equally weights the text embeddings associated with each text prompt.
@@
 [18], we alternate between the two text prompts.
-(iv) CLIP Min.: Score-based combination, denoted as CLIP-min, where we select the text prompt corresponding to the lowest CLIP score from the previous denoising iteration.
+(iv) CLIP Minimum: Score-based combination, denoted as CLIP-min, where we select the text prompt corresponding to the lowest CLIP score from the previous denoising iteration.
 Backbone architecture.
-We use the Stable Diffusion 2.1 backbone model in all our experiments. There is no training involved, we use the pretrained model to directly perform inference.
-Method CLIP-combined CLIP-add
-Vanilla $0.3463$ $0.3012$
-Linear Int. $0.2885$ $0.2778$
-Alt. Samp. $0.3445$ $0.3098$
-CLIP Min. $0.3195$ $0.2955$
-Step $0.322$ $0.2997$
-Black Scholes $0.3469$ $0.3112$
-Table 1: We follow prior work [9, 3] and analyze using CLIP Scores.
-The Black Scholes algorithm for prompt mixing in diffusion models achieves superior results, compared to other prompt-mixing techniques, as also evidenced by the qualitative results.
-5.1 Experimental Settings
+We use the Stable Diffusion 2.1 backbone model [30] in all our experiments.
+There is no training involved, we use the pretrained model to directly perform inference.
+Our experiments are run on one NVIDIA A5000 GPU with 24 GB memory.
+Method CLIP-combined ($\uparrow$) CLIP-add ($\uparrow$) BLIP $\odot$ DINO ($\uparrow$) KID ($\downarrow$) Steps ($\downarrow$) Time (s) ($\downarrow$) GPU hrs Memory (GB) ($\downarrow$)
+Linear Int. [17] $0.2885$ $0.2778$ $0.2588$ $0.02851$ $50$ $6.5$ $0.001805$ $7.1$
+Alt. Samp. [18] $0.3445$ $0.3098$ $0.3894$ $0.01786$ $100$ $14$ $0.00389$ $7.7$
+CLIP Min. $0.3195$ $0.2955$ $0.3107$ $0.00866$ $100$ $14$ $0.00389$ $7.7$
+Step [26] $0.3220$ $0.2997$ $0.3390$ $0.01709$ $100$ $14$ $0.00389$ $7.7$
+Black Scholes 0.3469 0.3112 0.3912 0.01531 100 14 0.00389 7.7
+Table 1: We evaluate various properties using CLIP Scores, BLIP Scores, DINO, and KID.
+These metrics help us assess overall text alignment with the combined text prompts, the preservation of attributes related to individual concepts, and the quality of the generated images.
+The Black Scholes algorithm for concept blending in diffusion models achieves superior results, compared to other prompt-mixing techniques, as also evidenced by the qualitative results.
+Hyperparameters.
+Based on our experiments for the vanilla combination using Stable Diffusion 2.1 for the dataset under consideration, where we found that a CLIP score of approximately 0.25 indicates reasonable text-image alignment, we opted for a constant value of 0.25 for the strike price.
+The ordering of prompts does not matter.
+This is because, at every step of diffusion denoising, the algorithm chooses the prompt that should be selected by computing the Black Scholes score wrt each prompt, and this process is agnostic to the ordering of the prompts.
+More details on hyperparameters and the backbone architecture can be found in the appendix.
+Figure 3: Our method can be extended to concept blending involving more than 2 prompts (3 in this case), as shown above.
+5.1 Analysis and Comparisons
+We construct a dataset with $4$ types of scenarios to analyze varying complexities, spanning:
+  • Simple text prompts: A single object per prompt.
+  • Multiple objects: Multiple objects described within one prompt.
+  • Object actions against backgrounds: Concept blending with respect to the object.
+  • Object performing action against background/style: Concept blending with respect to background and style.
+More details can be found in the appendix.
+Qualitative results are shown in Figure 1 and Figure 2, with additional examples provided in the appendix.
+Quantitative comparisons are summarized in Figure 1.
+A detailed analysis of our method relative to prior art is described below:
+  • Vanilla Stable Diffusion (SD): Vanilla Stable Diffusion v2.1 can generate plausible combinations of the input text prompts.
+However, due to hallucination issues and constraints imposed by the training distribution, it struggles to capture the distinct characteristics of individual prompts.
+For instance, in Figure 1 the unique features of the parrot, cat, dog/penguin, and sunset/penguin are not well preserved, and similarly, Figure 2 shows a lack of distinct characteristics for avocado/raccoon, muffin, parrot, and the oil painting style.
+  • Linear Interpolation [17]: Linear interpolation, which assumes a simple linear relationship, is ill-suited for blending two text prompts due to the highly non-linear nature of the text-image space.
+It only traverses a straight path between endpoints in the latent space, failing to capture nuanced variations or generate novel features.
+  • Alternating Sampling [18]: While alternating sampling produces features related to both prompts in the final output image, it often results in poorly defined objects, unrealistic appearance, and low overall quality.
+Artifacts are prevalent, and many images appear implausible, leading to a high KID score.
+The main issue is that the method alternates between the two text prompts without emphasizing the most relevant one during each denoising step.
+  • Step-wise Switching [18]: This technique suffers from similar issues as alternating sampling and fails to effectively capture the detailed characteristics of multiple text prompts.
+  • CLIP-min: The CLIP-min approach tends to favor one text prompt over the other, resulting in images that do not adequately reflect the intersection of both prompts.
+It selects the text prompt with the lowest CLIP score at each step, thereby overlooking the dynamic nature of the diffusion denoising process.
+  • Black-Scholes Algorithm: Our proposed method, based on the Black-Scholes algorithm, effectively generates realistic images that align with the intersection of the two text prompts.
+Our approach preserves the unique attributes of each prompt while minimizing unrealistic artifacts.
+By modeling the dynamics of the diffusion denoising process, our algorithm strategically selects the optimal text prompt at every step, achieving a balanced synthesis and superior quantitative performance.
+In summary, our Black-Scholes model–inspired method outperforms previous prompt mixing approaches by generating realistic images that meticulously preserve and balance the features of each individual text prompt.
+Computational complexity
+We report the computational requirements for different methods in Table 1.
+All values are computed using one NVIDIA A5000 GPU, on the 2 prompts blending scenario.
+Our algorithm incurs minimal computational overhead over other methods, primarily arising from the computation of the CLIP score of the generated image w.r.t.
+all individual text prompts at every step of the diffusion denoising based generation process.
+6 Conclusions, Limitations, and Future Work
+In this paper, we introduced a novel approach to concept blending for text-to-image diffusion models that is deeply inspired by financial probabilistic models, particularly the Black-Scholes framework.
+By drawing an analogy between asset pricing and the dynamics of the diffusion process, we developed an automated algorithm that dynamically selects the most relevant text prompt at each denoising step.
+This approach not only enhances the alignment between the generated image and the input prompts but also offers a data-efficient alternative to conventional prompt mixing methods.
+Our extensive qualitative and quantitative evaluations demonstrate that our method outperforms traditional techniques such as linear interpolation, alternating sampling, and manual prompt switching.
+While our approach leverages the strengths of modern diffusion models, it is important to note that it may not extend to non-Gaussian diffusion processes [6] or one-step diffusion models [39]—a limitation that opens interesting avenues for future research.
+Moreover, the underlying principles of our method have significant implications beyond concept blending.
+They can be readily integrated into state-of-the-art text-to-image architectures, and they pave the way for critical downstream tasks including image editing [17, 37, 4], compositional synthesis [24, 3], handling complex prompts [23], text-based view synthesis [18, 19], and personalized image generation [31, 20].
+Our work demonstrates that financially inspired methods have the potential to significantly enhance concept blending in text-to-image generation, thereby opening avenues for broader innovation across the field of generative AI.
+Impact Statement.
+Our approach advances the state-of-the-art in generative AI for content creation by harnessing the synergy between diffusion models and financial probabilistic frameworks.
+However, the increased realism and flexibility afforded by our method also raise significant concerns regarding potential misuse.
+This dual-edged potential underscores the urgent need for continued research in watermarking and deepfake detection to responsibly mitigate associated risks while enabling creative innovations.
+Acknowledgements.
+This work was supported in part by ARO Grants W911NF2110026, W911NF2310046, W911NF2310352 and Army Cooperative Agreement W911NF2120076.
+A.1. Experimental Settings
 We consider the following experimental settings:
@@
 Set 3:
-Object actions against backgrounds (prompt mixing w.r.t.
+Object actions against backgrounds (concept blending w.r.t.
 object).
@@
 Set 4:
-Object performing action against a background (prompt mixing w.r.t.
+Object performing action against a background (concept blending w.r.t.
 background).
@@
 Overall, we have $90$ prompts in this experiment.
-5.2 Analysis and Comparisons
-We show qualitative results in Figure 1 and Figure 2.
-More qualitative results can be found in the appendix.
-The quantitative comparisons are in Figure 1.
-Our quantitative results are averaged across the metrics for each of the four sets, with $680$, $355$, $1050$, $450$ images respectively.
-A detailed analysis of our method and benefits over prior art is as follows:
-  • Vanilla stable diffusion (SD): Vanilla stable diffusion v2.1 can create plausible combinations of the provided text prompts.
-However, due to hallucination issues, it is constrained by the distributions it learned during training.
-As a result, it struggles to capture clear characteristics of individual text prompts.
-For example, in Figure 1, the distinct features of the parrot, cat, dog/penguin, and sunset/penguin are missing.
-Similarly, in Figure 2, the avocado/raccoon, muffin, parrot, and oil painting style lack distinct characteristics.
-Despite this, the generated images align with the broad text prompt descriptions, resulting in a considerable CLIP score.
-  • Linear interpolation [17]: Linear interpolation is ill-suited for generating an image that combines two text prompts.
-This limitation arises from the non-linear nature of the text-image space.
-While linear interpolation assumes a simple linear relationship, the actual mapping between textual descriptions and image features is intricate.
-Furthermore, linear interpolation can only traverse a straight path between two points in the latent space, failing to capture nuanced variations or produce novel features beyond the endpoints.
-  • Alternating Sampling [18],: Alternating sampling generates features related to both text prompts in the final output image.
-However, the objects produced suffer from poor definition, unrealistic appearance, and low quality in many instances.
-Artifacts are prevalent, and several generated images appear implausible.
-The underlying issue lies in the algorithm’s routine alternation between the two text prompts during diffusion denoising, without adequately emphasizing the most relevant prompt at each step.
-Despite these limitations, the model achieves a reasonable CLIP Score by broadly aligning with the overall image description, akin to vanilla denoising.
-  • Step-wise switching [18]: Step-wise switching is similarly ineffective for the same reasons as Alternating sampling.
-  • CLIP-min: The CLIP-min results exhibit bias toward one of the text prompts in most cases, preventing the model from generating an image that aligns with the intersection of both prompts.
-This bias arises because the model overlooks the dynamics of diffusion denoising and instead selects the text prompt with the lowest CLIP score at each step, resulting in bias-related issues.
-  • Black Scholes algorithm: Our approach based on the Black-Scholes algorithm effectively generates realistic images that align with the intersection of the two text prompts.
-These images exhibit minimal unrealistic artifacts.
-Additionally, our algorithm successfully preserves individual characteristics corresponding to each text prompt.
-By modeling the dynamics of diffusion denoising, the algorithm strategically selects the optimal text prompt at each step, achieving a balanced synthesis of both prompts.
-Notably, our model attains superior quantitative results.
-In summary, the Black-Scholes model outperforms all previous methods for prompt mixing by generating realistic images that meticulously preserve and balance the characteristics of each individual text prompt.
-Figure 3: Image variations from our method using the Black Scholes model for a single prompt, starting from different random Gaussian noise initializations.
-6 Conclusions, Limitations, and Future Work
-In this paper, we introduce a novel perspective on prompt mixing, inspired by a probablistic prediction model and its theoretical framework from the domain of Finance.
-We draw parallels between the Black-Scholes algorithm used in financial markets and diffusion models.
-From this comparison, we derive an algorithm for prompt mixing, aiming to generate images that align with multiple text prompts.
-Our extensive analysis showcase both qualitative and quantitative advantages of our approach.
-Our study has a few limitations, which opens avenues for future work: (i) We follow prior work [9, 3] by using CLIP scores for quantitative evaluation.
-However, CLIP scores may not be the optimal evaluation strategy, as they can yield similar scores for images with significant quality differences.
-Due to its learned nature, CLIP is vulnerable to hallucination and bias issues related to its training data.
-Consequently, it may assign similar scores to images that exhibit general characteristics of the text prompt, without focusing on specific details and features of the scene entities described in the text prompt.
-To address this, future research could explore more advanced image-language models [25] for evaluation; the development of good metrics for the evaluation of generative models is a widespread issue, and could help better quantify the performance of our method.
-(ii) Our study focuses on prompt mixing with two prompts.
-Investigating the impact of using more than two prompts could be a promising direction for future work.
-(iii) While our method leverages diffusion models, it may not be applicable to non-Gaussian [6] diffusion models or one-step diffusion models [41], as demonstrated in recent research.
-Moreover, future work on integrating our approach with recent advancements in attention guidance, layout modeling, and related areas could extend its applicability to critical downstream tasks such as image editing [17, 39, 4], compositionality [24, 3], handling complex prompts [23], text-based view synthesis [18, 19] and personalized image generation [32, 20].
-While recognizing the success of recent text-to-image models like Dalle-3, Stable Diffusion 3, Firefly 3, and Imagen 3 in generating images from diverse and complex text prompts, we introduce a data-efficient approach to prompt mixing.
-Our method offers advantages over previous techniques for prompt-mixing, and future extension could explore integrating our model’s benefits into state-of-the-art text-to-image architectures for data efficient prompt mixing and downstream applications.
-Societal impact.
-While our method offers valuable tools for generative AI-based content creation, its potential misuse underscores the need for research in watermarking and deepfake detection to mitigate risks.
-Acknowledgements.
-This work was supported in part by ARO Grants W911NF2110026, W911NF2310046, W911NF2310352 and Army Cooperative Agreement W911NF2120076.
+Figure 4: Image variations from our method using the Black Scholes model, starting from different random Gaussian noise initializations.
+Figure 5: Ablation experiment on the strike price $K$ reveals that the results are optimal at values of $K$ close to the CLIP score of the image generated in the vanilla case using a combination of the constituent text prompts.
+A.1.1. Hyperparameters:
+We use the Stable Diffusion 2.1 backbone in all our experiments.
+We use standard diffusion parameters: image size - (512, 512); classifier-free guidance scale - 7.5; DDIMScheduler with betaStart set to 0.00085, betaEnd set to 0.012, betaSchedule set to “scaledLinear”, clipSample set to False and setAlphaToOne set to False.
+We use $100$ inference steps in all our experiments.
+For the set of prompts in our dataset under consideration, the average CLIP score for the combination of prompts in the vanilla case was $0.25$. Hence, we set the value of strike price $K$ at $0.25$.
+A.2. Assumptions made by the Black Scholes Algorithm
+The Black Scholes model makes five assumptions: (i) No dividends are paid out during the life of an option, (ii) Market movements are somewhat random, (iii) There are no transaction costs in buying the asset, (iv) The volatility and risk free rate of the underlying asset are known and constant, (v) The returns of the underlying asset are normally distributed.
+The relation between the assumptions made by the Black Scholes algorithm and diffusion models is as follows:
+  • No Dividends (Assumption i): This assumption is specific to financial options and does not directly apply to diffusion models.
+In the context of diffusion models, we can consider that the "output" (the generated image) does not have any intermediate rewards or returns, similar to how options do not yield dividends.
+  • Random Market Movements (Assumption ii): This assumption aligns well with the inherent randomness in the diffusion process.
+Just as stock prices are subject to unpredictable fluctuations, the generation of images in diffusion models involves stochastic processes where the final output is influenced by random noise and iterative refinement.
+  • No Transaction Costs (Assumption iii): Similar to the first assumption, transaction costs are not relevant in the context of diffusion models.
+Instead, we can think of the computational resources and time required for generating images as analogous to transaction costs, but they do not affect the model’s fundamental operation.
+  • Constant Volatility and Risk-Free Rate (Assumption iv): In diffusion models, while volatility is not explicitly defined as in financial contexts, the concept of stability in the generation process can be likened to having a consistent framework for how noise is added and refined.
+The “risk-free rate” does not have a direct counterpart, but the idea of a stable environment for generating images can be considered.
+  • Normally Distributed Returns (Assumption v): This assumption can be related to the distribution of pixel values or features in the generated images.
+In many diffusion models, the outputs are conditioned on a distribution that can be approximated as normal, especially when considering the latent space representations.
+In summary, while some assumptions of the Black-Scholes model are not directly applicable to diffusion models, others, particularly the randomness and distribution aspects, provide a meaningful connection, which leads to our algorithm on how concepts from finance can inform generative modeling techniques!
+Figure 6: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts.
+For instance, it misses out on the characteristics of the dog, parrot and ice-cream.
+Linear interpolation generates images that are not consistent with the text prompts.
+CLIP Min.
+generates images biased towards one of the text promots.
+Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases.
+For instance, in order, the issues are: missing characteristics of dog, missing characteristics of coffee mug and artifacts in teddy bear, artifacts and missing characteristics of parrot, missing characteristics of teddy bear and artifacts in teddy bear, missing characteristics of rabbit.
+Our Black Scholes method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts.
+The text prompts are from set 1.
+Figure 7: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts.
+For instance, it misses out on the characteristics of the dog, cat, corgi dog and basket/ teapot mixing.
+Linear interpolation generates images that are not consistent with the text prompts.
+CLIP Min.
+is biased towards one of the text prompts.
+Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases.
+For instance, in order, the issues are: artifacts in dog/banana and missing characteristics of dog, missing characteristics of cat, the corgi dog in the dog/apple image does not describe the dog as well as our Black Scholes method does, the dog is not well described in the fourth example, the teapot is not well described in the final image.
+Our method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts.
+The text prompts are from set 2.
+Figure 8: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts.
+For instance, in the first and fifth image, the characteristics of the teddy bear are missing, the second, third and fourth image do not describe the cat well, Linear interpolation generates images that are not consistent with the text prompts.
+CLIP Min is again biased towards one of the text prompts.
+Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases.
+Specifically, the first image does not describe the subjects well, the second and third images are reasonably generated but our Black Scholes method generates a clearer image of the penguin blended cat, and teddy bear blended cat, the results of Alt.
+for the fourth case has artifacts and Step does not describe the cat well, the teddy bear is not well described in the fifth image.
+Our Black Scholes method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts.
+The text prompts are from set 3.
+Figure 9: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts.
+The first two images are cartoon-ish, and Northern lights is not clearly depicted.
+The third image also misses a clear description of Northern lights.
+In the fourth image, some of the billboards in Times Square are green, but there is no sign of Northern lights.
+In the fifth image, the sunset is missing.
+Linear interpolation generates images that are not consistent with the text prompts.
+CLIP Min results are biased towards one of the prompts for all images, except the second one.
+Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases.
+In the first image, the kangaroo is not generated well (almost flying in Alt’s result, and the characteristics of the kangaroo are not well generated by Step).
+In the second image, the characteristics of the tulip garden are not well represented.
+In the third image, notice how our model generates the characteristics of the sunset and Northern lights better, while ensuring that they are blended well to form a realistic image.
+In the fourth image, Alt and Step miss out on the characteristics of Times Square and Northern lights respectively.
+In the fifth image, Alt and Step miss out on the characteristics of Times Square and the sunset respectively.
+Our Black Scholes method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts.
+The text prompts are from set 4.